OpenStudy (anonymous):

(5x+5)/(x^2+8x+15)=0

OpenStudy (anonymous):

i started factoring and got 5(x+1)/(x+5)(x+3)=0 but im not sure what to do after....so i just take the denominator and try to figure it out from there?

OpenStudy (anonymous):

Do you have to find restrictions and/or solve for x?

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

Alright. You would have the right approach if you were simplifying. However, here, just solve for x regularly. \[\frac{5x + 5}{x^2 + 8x + 15} = 0 \implies 5x + 5 = 0 \times (x^2 + 8x = 15) \implies 5x + 5 = 0\]Can you solve from there?

OpenStudy (anonymous):

im not sure if i understand what you did

OpenStudy (anonymous):

Alright. You have x in both the denominator and numerator. Therefore, you multiply both sides by what's in the denominator so that you can solve for x.

OpenStudy (anonymous):

Since 0 times anything is 0, that cancels out.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so now you just solve for x here: 5x + 5 = 0

OpenStudy (anonymous):

super easy! i get it

OpenStudy (anonymous):

Good :)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

np :D

OpenStudy (anonymous):

how would you do it if its like this

OpenStudy (anonymous):

(x-7)/5 + 1/5 = -x/10